OpenStudy (calculusxy):

MEDAL for Scientific Notation Operations help!!! 5.4 x 10^2 x 7.5 x 10^{-7}

OpenStudy (calculusxy):

\[\huge 5.4 \times 10^2 \times 7.5 \times 10^{-7}\]

imqwerty (imqwerty):

scientific notation of 23.212 x10^3 is 2.3212 x 10^4

imqwerty (imqwerty):

jst an example^

OpenStudy (calculusxy):

\[\large (5.4 \times 7.5)(10^2 \times 10^{-7}) = 40.50 \times 10^{-5}\] I don't know how to go from there.

OpenStudy (welshfella):

first find 5.4 * 7.5 then 10^2 * 10^-7

OpenStudy (welshfella):

right now that's not scientific notation because 40.5 is not between 1 and 10

OpenStudy (welshfella):

Hint: part of the next procedure is to divide 40.5 by 10

OpenStudy (calculusxy):

okay. so that would be 4.05

OpenStudy (welshfella):

- but os corse you have to retain the correct value of the number

OpenStudy (calculusxy):

\[4.05 \times 10^{-1} \times 10^{-5}\]

OpenStudy (calculusxy):

would it be something like that?

OpenStudy (welshfella):

you have to multiply the exponent part by 10

OpenStudy (welshfella):

you are not correct in the last post

OpenStudy (calculusxy):

okay. so can you explain to me how i should do it?

OpenStudy (welshfella):

its 4.05 * 10^1 * 10^-5

OpenStudy (welshfella):

since we divided 40.5 by 10 we need to do the inverse, multiply, exponent part by 10 to reatain the correct value

OpenStudy (calculusxy):

oh is it because i have to multiply 10^1 to get to 40.50 and then go about from there?

OpenStudy (calculusxy):

oh ok. so the answer would be \[\large 4.5^{-4}\]

OpenStudy (welshfella):

No its 4.05 * 10^-4

OpenStudy (calculusxy):

oh yeah sorry typo

OpenStudy (welshfella):